b. xj(t) = xi(t) – 2x2(t); Xz(t) = 2x1(t) + x2(t); x1(0) = 0, x2(0) = 4 2
Answer #1

Ⓡ X(t)= x1(+)-2xy (t) i 400) = 0 (t) = 2.2;(+) + X2 () · XCO)= 4. Matrix form of the given system is. x (t)= A XV X(0)=b (x= et е et (Cos (24) + i sin(24) ) [ 1 ] Seuleris formula} I cos (24) - sin (24) {multiply cosines of Cos (24) tisin (at). si